# PuzzleBench: Can LLMs Solve Challenging First-Order Combinatorial Reasoning Problems?

### v2

@@ Appendix B Prompt Templates
In this section we provide prompt templates used for our experiments on PuzzleBench, including the templates for the baselines we experimented with, Puzzle-LM as well as prompt templates for providing feedback.
@@ B.1 Few-Shot Prompt Template
Task: <Description of the Rules of the problems> Input-Format: <Description of Textual Representation of Inputs> <Input Few Shot Example-1> <Input Few Shot Example-2> ........................ ........................ <Input Few Shot Example-n> Output-Format <Description of Textual Representation of Outputs> <Output of Few Shot Example-1> <Output of Few Shot Example-2> ............................ ............................ <Output of Few Shot Example-n> Input problem instance to be solved: <Puzzle from the Test Set>
@@ B.2 PAL Prompt Template
The following box describes the base prompt template used for PAL experiments with PuzzleBench. Write a Python program to solve the following problem: Task: <Description of the Rules of the problem> Input-Format: <Description of Textual Representation of Inputs> Sample-Input: <Sample Input from Feedback Set> Output-Format: <Description of Textual Representation of Outputs> Sample-Output: <Output of Sample Input from Feedback Set> Don’t write anything apart from the Python program; use Python comments if needed. The Python program is expected to read the input from input.txt and write the output to a file named output.txt. The Python program must only use standard Python libraries.
B.3 Puzzle-LM Template
@@ B.3.1 Base Prompt
Write a Python program to solve the following problem: Task: <Description of the Rules of the problem> Input-Format: <Description of Textual Representation of Inputs> Sample-Input: <Sample Input from Feedback Set> Output-Format: <Description of Textual Representation of Outputs> Sample-Output: <Output of Sample Input from Feedback Set> The Python program must read the input from input.txt and convert that particular input to the corresponding constraints, which it should pass to the Z3 solver, and then it should use the Z3 solver’s output to write the solution to a file named output.txt Don’t write anything apart from the Python program; use Python comments if needed.
@@ B.4 Feedback Prompt Templates
These prompt templates are used to provide feedback in the case of Puzzle-LM or PAL.
@@ B.4.2 Verification Error
Your code is incorrect, when run on the input: <INPUT> the output produced is <OUTPUT-GENERATED> which is incorrect whereas one of the correct output is <GOLD-OUTPUT>. Rewrite your code and fix the mistake.
@@ B.4.3 Timeout Error
Your code was inefficient and took more than <TIME-LIMIT> seconds to execute for the following input: <INPUT>. Rewrite the code and optimize it.
@@ B.5 Logic-LM Prompt Template
The following box describes the prompt for Logic-LM experiments with PuzzleBench, the prompt is used to convert the input to its symbolic representation. Task: <Description of the Rules of the problem> Input-Format: <Description of Textual Representation of Inputs> Sample-Input: <Sample Input from Feedback Set> Output-Format: <Description of Textual Representation of Outputs> Sample-Output: <Output of Sample Input from Feedback Set> Input problem to be solved: <Puzzle from the Test Set> The task is to declare variables and the corresponding constraints on them in SMT2 for the input mentioned above. The variables and constraints should be such that once the variables are solved for, one can use the solution to the variables (which satisfies the constraints) to get to the output in the desired format for the above mentioned input. Only Write the SMT2 code and nothing else. Write the complete set of SMT2 variables and constraints. Enclose SMT2 code in ‘‘‘smt2 ‘‘‘
@@ Appendix C Experimental Details
C.1 PuzzleBench
All methods are evaluated zero-shot, meaning no in-context demonstrations for the task are provided to the LLM. We choose the zero-shot setting for PuzzleBench because of the structured nature of problems, making it unfair to provide demonstrations of highly related problems instances to the LLM. The LLM is only given a description of the rules of the problem and the task it has to perform. For PAL and Puzzle-LM we present results with 10 solved examples for feedback.
C.2 Hardware Details
@@ D.1 Inference Time
The following tables describes the average inference time for test set instances of a few illustrative problems in PuzzleBench.
Problem PAL Puzzle-LM
Sudoku
2.01 0.215
Latin Square
5.46 0.2
Cryptarithmetic
0.83 0.73
Independent Set
1.438 0.106
Minimum Dominating Set
0.98 0.112
Sujiko
0.742 0.102
Vertex Cover
1.58 0.105
Table 6: Average inference time in seconds of Puzzle-LM and PAL for test set instances for selected problems in PuzzleBench
Puzzle-LM performs much better compared to PAL because PAL programs often tend to be brute force and inefficient whereas the solver can exploit the nature of the input-instance while performing the reasoning with Puzzle-LM.
@@ Appendix E Examples
E.1 Puzzle-LM
This section includes example programs generated by Puzzle-LM for some illustrative problems in PuzzleBench. Each program reads the input from a file, generates the corresponding constraints, calls the solver internally and then uses the solution from the solver to write the output in the desired format to a file.
Figure 12: Puzzle-LM example: correct program for sudoku generated by GPT-4-Turbo.
Figure 13: Puzzle-LM example: correct program for keisuke generated by GPT-4-Turbo.
Figure 14: Puzzle-LM example: correct program for hamiltonian path generated by GPT-4-Turbo.
Figure 15: Puzzle-LM example: correct program for vertex-cover generated by GPT-4-Turbo.
Figure 16: Puzzle-LM example: snippet of incorrect program for binairo generated by GPT-4-Turbo and same snippet after correction by feedback.
Figure 17: Puzzle-LM example: correct program for subset-sum generated by GPT-4-Turbo.
@@ E.2 PAL
This section includes example programs generated by PAL for some illustrative problems in PuzzleBench. Each program reads the input from a file, performs the reasoning and writes the output to another text file.
@@ Appendix F Logic-LM
This section describes example runs of Logic-LM for certain problems in PuzzleBench.

### v3

@@ Appendix B Prompt Templates
In this section we provide prompt templates used for our experiments on FCoReBench, including the templates for the baselines we experimented with, SymPro-LM as well as prompt templates for providing feedback.
@@ B.1 Few-Shot Prompt Template
{mdframed}
Task:
<Description of the Rules of the problems>
Input-Format:
<Description of Textual Representation of Inputs>
<Input Few Shot Example-1>
<Input Few Shot Example-2>
<Input Few Shot Example-n>
Output-Format
<Description of Textual Representation of Outputs>
<Output of Few Shot Example-1>
<Output of Few Shot Example-2>
<Output of Few Shot Example-n>
Input problem instance to be solved:
<Problem Instance from the Test Set>
@@ B.2 PAL Prompt Template
The following box describes the base prompt template used for PAL experiments with FCoReBench. {mdframed} Write a Python program to solve the following problem:
Task:
<Description of the Rules of the problem>
Input-Format:
<Description of Textual Representation of Inputs>
Sample-Input:
<Sample Input from Feedback Set>
Output-Format:
<Description of Textual Representation of Outputs>
Sample-Output:
<Output of Sample Input from Feedback Set>
Don’t write anything apart from the Python program; use Python comments if needed.
The Python program is expected to read the input from input.txt and write the output to a file named output.txt.
The Python program must only use standard Python libraries.
B.3 SymPro-LM Template
@@ B.3.1 Base Prompt
{mdframed}
Write a Python program to solve the following problem:
Task:
<Description of the Rules of the problem>
Input-Format:
<Description of Textual Representation of Inputs>
Sample-Input:
<Sample Input from Feedback Set>
Output-Format:
<Description of Textual Representation of Outputs>
Sample-Output:
<Output of Sample Input from Feedback Set>
The Python program must read the input from input.txt and convert that particular input to the corresponding constraints, which it should pass to the Z3 solver, and then it should use the Z3 solver’s output to write the solution to a file named output.txt
Don’t write anything apart from the Python program; use Python comments if needed.
@@ B.4 Feedback Prompt Templates
These prompt templates are used to provide feedback in the case of SymPro-LM or PAL.
@@ B.4.2 Verification Error
Your code is incorrect, when run on the input: <INPUT> the output produced is <OUTPUT-GENERATED> which is incorrect whereas one of the correct output is <GOLD-OUTPUT>.
Rewrite your code and fix the mistake.
@@ B.4.3 Timeout Error
Your code was inefficient and took more than <TIME-LIMIT> seconds to execute for the following input: <INPUT>.
Rewrite the code and optimize it.
@@ B.5 Logic-LM Prompt Template
The following box describes the prompt for Logic-LM experiments with FCoReBench, the prompt is used to convert the input to its symbolic representation. {mdframed} Task:
<Description of the Rules of the problem>
Input-Format:
<Description of Textual Representation of Inputs>
Sample-Input:
<Sample Input from Feedback Set>
Output-Format:
<Description of Textual Representation of Outputs>
Sample-Output:
<Output of Sample Input from Feedback Set>
Input problem to be solved:
<Problem Instance from the Test Set>
The task is to declare variables and the corresponding constraints on them in SMT2 for the input mentioned above. The variables and constraints should be such that once the variables are solved for, one can use the solution to the variables (which satisfies the constraints) to get to the output in the desired format for the above mentioned input.
Only Write the SMT2 code and nothing else. Write the complete set of SMT2 variables and constraints. Enclose SMT2 code in ‘‘‘smt2 ‘‘‘
B.6 ToT
In this section we give an example of the ToT prompts used for experiments on FCoReBench. We use latin square as the running example.
B.6.1 Propose Prompt
This prompt is called for each search node to get the possible next states. {mdframed} Task:
We are given a nxn partially solved board and have to solve it according to the following rules:
- We need to replace the 0s with numbers from 1-n.
- Non-zero numbers on the board cannot be replaced.
- Each number from 1-n must appear exactly once in each column and row in the solved board
Given a board, decide which cell to fill in next and the number to fill it with, each possible next step is separated by a new line.
You can output up-to 3 next steps.
If the input board is fully filled or no valid next step exists output only ’END’.
Sample-Input-1:
1 0 3
2 0 0
0 1 2
Possible next steps for Sample Input-1:
1 2 3
2 0 0
0 1 2
1 0 3
2 0 0
3 1 2
1 0 3
2 3 0
0 1 2
Sample-Input-2:
1 2 3
2 3 1
3 1 2
Possible next steps for Sample Input-2:
END
Input:
<node from the search tree>
Possible next steps for Input:
B.6.2 Value Prompt
This prompt is called for each search node to evaluate how likely it is to get to the solution from that node. We use this to prune the search tree. {mdframed} Task:
We are given a nxn partially solved board and have to solve it according to the following rules:
- We need to replace the 0s with numbers from 1-n.
- Non-zero numbers already on the board cannot be replaced.
- Each number from 1-n must appear exactly once in each column and row in the solved board.
Given a partially filled board, evaluate how likely it is to reach a valid solution (sure/likely/impossible)
Output-Format:
The output should have two lines as follows:
<Reasoning>
<Sure/Likely/Impossible>
Sample-Input-1:
0 0 0
0 0 0
0 0 0
Board is empty, hence it is always possible to get to a solution.
Sure
Sample-Input-2:
1 0 3
2 0 0
0 1 2
No constraint is violated till now and it is likely to get to a solution.
Likely
Sample-Input-3:
1 1 3
2 0 0
0 1 2
Constraint violated in first row.
Impossible
Input:
<node from the search tree>
@@ Appendix C Experimental Details
C.1 FCoReBench
All methods are evaluated zero-shot, meaning no in-context demonstrations for the task are provided to the LLM. We choose the zero-shot setting for FCoReBench because of the structured nature of problems, making it unfair to provide demonstrations of highly related problems instances to the LLM. The LLM is only given a description of the rules of the problem and the task it has to perform. For PAL and SymPro-LM we present results with 10 solved examples for feedback.
C.1.1 ToT prompting
We evaluate ToT prompting [Yao et al., 2023] on 3 problems in FCoReBench. Our implementation closely resembles the official implementation which we adapt for grid based logical puzzles. We use a BFS based approach with propose and value prompts. An example prompt for latin square can be found in Appendix B.6. Problems in our benchmark have huge branching factors, to reduce experimentation cost, we greedily prune the search frontier to 5 nodes at each depth based on scores assigned by the LLM during the value stage. Additionally during the propose stage we prompt the LLM to output at most 3 possible next steps. The temperature is set to 0.0 for reproducibility. Unlike the original implementation problems in our benchmark can have varying search depths, hence we explicitly ask the LLM to output ’END’ once a terminal node is reached. At any depth if a terminal node is amongst the best nodes we terminate the search and return the terminal nodes at that depth, otherwise we search till a maximum search depth governed by the problem instance.
C.2 Other Datasets
We evaluate SymPro-LM on 3 other datasets apart from FCoReBench. Our evaluation closely follows Logic-LM’s evaluation [Pan et al., 2023]. For baselines we use the same prompts as Logic-LM. Logic-LM did not evaluate PAL, for which we write prompts on our own similar to the CoT prompts used by Logic-LM. For SymPro-LM we write prompts on our own. We use the same in-context examples as used for Logic-LM. We instruct the LLM to write a Python program to parse the input problem, setup variables/constraints and pass these to a symbolic solver, call the solver and using the solver’s output print the final answer. For LogicalDeduction we use the python-constraints  [^3] https://github.com/python-constraint/python-constraint package which is a CSP solver. For other datasets we use the Z3-solver  [^4] https://pypi.org/project/z3-solver/. Since all problems are single correct MCQ questions we use accuracy as our metric. Like Logic-LM if there is an error during execution of the program generated by the LLM we fall back on using chain-of-thought to predict the answer. The following sections provide descriptions for the datasets used.
C.2.1 PrOntoQA
PrOntoQA [Saparov and He, 2023a] is a recent synthetic dataset created to analyze the deductive reasoning capacity of LLMs. We use the hardest fictional characters version of the dataset, based on the results in [Saparov and He, 2023a]. Each version is divided into different subsets depending on the number of reasoning hops required. We use the hardest 5-hop subset for evaluation. Each question in PrOntoQA aims to validate a new fact’s veracity, such as “True or false: Alex is not shy.” The following box provides an example: {mdframed} Context: Each jompus is fruity. Every jompus is a wumpus. Every wumpus is not transparent. Wumpuses are tumpuses. Tumpuses are mean. Tumpuses are vumpuses. Every vumpus is cold. Each vumpus is a yumpus. Yumpuses are orange. Yumpuses are numpuses. Numpuses are dull. Each numpus is a dumpus. Every dumpus is not shy. Impuses are shy. Dumpuses are rompuses. Each rompus is liquid. Rompuses are zumpuses. Alex is a tumpus
Question: True or false: Alex is not shy.
Options:
A) True
B) False
C.2.2 ProofWriter
[…]
Question: Based on the above information, is the following statement true, false, or unknown? Anne is white.
Options:
A) True
B) False
C) Uncertain
C.2.3 LogicalDeduction
LogicalDeduction bench authors, 2023 is a challenging logical reasoning task from the BigBench collaborative benchmark. The problems are mostly about deducing the order of a sequence of objects from a minimal set of conditions. We use the full test set consisting of 300 examples. The following box provides an example: {mdframed} Context: The following paragraphs each describe a set of three objects arranged in a fixed order. The statements are logically consistent within each paragraph. In an antique car show, there are three vehicles: a station wagon, a convertible, and a minivan. The station wagon is the oldest. The minivan is newer than the convertible.
Question: Which of the following is true?
Options:
A) The station wagon is the second-newest.
B) The convertible is the second-newest.
C) The minivan is the second-newest.
C.3 Hardware Details
@@ D.1 Inference Time
The following tables describes the average inference time for test set instances of a few illustrative problems in FCoReBench.
Problem PAL SymPro-LM
Sudoku 2.01 0.215
Latin Square 5.46 0.2
Cryptarithmetic 0.83 0.73
Independent Set 1.438 0.106
Minimum Dominating Set 0.98 0.112
Sujiko 0.742 0.102
Vertex Cover 1.58 0.105
Table 8: Average inference time in seconds of SymPro-LM and PAL for test set instances for selected problems in FCoReBench
SymPro-LM performs much better compared to PAL because PAL programs often tend to be brute force and inefficient whereas the solver can exploit the nature of the input-instance while performing the reasoning with SymPro-LM.
@@ Appendix E Examples
E.1 SymPro-LM
E.1.1 FCoReBench
This section includes example programs generated by SymPro-LM for some illustrative problems in FCoReBench. Each program reads the input from a file, generates the corresponding constraints, calls the solver internally and then uses the solution from the solver to write the output in the desired format to a file.
Figure 13: SymPro-LM example: correct program for sudoku generated by GPT-4-Turbo.
Figure 14: SymPro-LM example: correct program for keisuke generated by GPT-4-Turbo.
Figure 15: SymPro-LM example: correct program for hamiltonian path generated by GPT-4-Turbo.
Figure 16: SymPro-LM example: correct program for vertex-cover generated by GPT-4-Turbo.
Figure 17: SymPro-LM example: snippet of incorrect program for binairo generated by GPT-4-Turbo and same snippet after correction by feedback.
Figure 18: SymPro-LM example: correct program for subset-sum generated by GPT-4-Turbo.
E.1.2 Other Datasets
Figure 19: SymPro-LM PrOntaQA Example Program.
Figure 20: SymPro-LM ProofWriter Example Program.
Figure 21: SymPro-LM LogicalDeduction Example Program.
@@ E.2 PAL
This section includes example programs generated by PAL for some illustrative problems in FCoReBench. Each program reads the input from a file, performs the reasoning and writes the output to another text file.
@@ Appendix F Logic-LM
This section describes example runs of Logic-LM for certain problems in FCoReBench.
@@
Figure 31: Logic-LM example: errors (highlighted comments) in constraints for sudoku instance generated by GPT-4-Turbo.